OpenStudy (anonymous):

You have a loaded coin that comes up heads with a 0.6 probability. Now you add a fair coin (probability heads = 0.5). You now have two coins - one loaded and one fair. You select a coin at random with a probability of 0.5, you flip it twice and it comes up heads twice. What is the probability that you picked the loaded coin?

OpenStudy (anonymous):

Hint: follow conditional probability theorem:

OpenStudy (anonymous):

The probability that the following is the answer is 0.6 :\[\frac{\frac{1}{2}\left(\frac{6}{10}\right)^2}{\frac{1}{2}\left(\frac{6}{10}\right)^2+\frac{1}{2}\left(\frac{1}{2}\right)^2}=\frac{36}{61} \]

OpenStudy (anonymous):

You have two probabilities: In coin 1 The probability that two heads come up. = 1/4 = 25/100 In coin 2 The probability that two heads come up. = 9/25 = 36/100 so you can conclude that: 9/25 > 1/4 You pick either randomly 50-50 What you want to know that according to the evidence that 2 heads came up from your chosen coin what is the likeliness that this chosen coin is the 9/25 coin. You can tell by simple logic that as 9/25 > 1/4, the loaded coin is more likely to be the one you chose as it obeys the probabilities more than the other coin. The exact number I don't know how to get it, I speculate it must be some sort of sum or difference between these two numbers ( 9/25, 1/4) maybe 9/25+1/4 = 61/100 from here sincerely I don't know, I have never heard of such problem on probabilities where you use posterior evidence to calculate previous decisions